OpenStudy (anonymous):

For children between ages 6 and 10, height y (in inches) is frequently a linear function of age t (in years). The height of a certain child is 45 inches at age 6 and 47.5 inches at age 7. (a) Express y as a function of t.

OpenStudy (anonymous):

y=t+0.5?

OpenStudy (campbell_st):

so you have 2 data points (7, 47.5) and (6, 45) do you know how to find the slope from 2 ordered pairs..?

OpenStudy (anonymous):

you have 2 points (6,46), (7,47.5) it says its linear so you want to find equation of line y=mx+b is standard form for line slope = (y1-y2)/(x1-x2) slope = (47.5-46)/(7-6) = 1.5/1 y= 1.5x +b 46 = 1.5(6) +b 46= 9 + b b = 37 eq: y = 1.5x +37

OpenStudy (campbell_st):

slight calculation error the point you've used should be (6, 45) not (6, 46)

OpenStudy (anonymous):

o right sorry my mistake slope should be (47.5 - 45)/(7-6) = 2.5 y = 2.5 + b 45 = 2.5(6) + b b = 30 eq: y= 2.5x + 30

OpenStudy (anonymous):

tried that. didn't work

OpenStudy (anonymous):

2.5t+30 was the answer